# VidMuse: A Simple Video-to-Music Generation Framework with Long-Short-Term Modeling

### v1

@@ 3.4 Evaluation Metrics
To quantitatively evaluate the effectiveness of our model, we employ a series of metrics to assess different models in terms of quality, fidelity, and diversity for generated music. Besides, we also examine the alignment degree between the video and generated music. We believe these metrics can thoroughly evaluate the performance of different methods.
Frechet Audio Distance (FAD) [^1] https://github.com/haoheliu/audioldm_eval [38] is a reference-free evaluation metric for assessing audio quality. Similar to Frechet Image Distance (FID)[32], it compares the embedding statistics of the generated audio clip with groundtruth audio clips. A shorter distance usually denotes better human-perceived acoustic-level audio quality. However, this metric cannot reflect semantic-level information in audio. We report the FAD based on the VGGish[31] feature extractor.
Frechet Distance (FD).33footnotemark: 3 measures the similarity between generated samples and target samples in audio generation fields. It’s similar to FAD but uses a PANNs feature extractor instead. PANNs[39] have been pre-trained on AudioSet[24], one of the largest audio understanding datasets, thus resulting in a more robust metric than FAD.
Kullback-Leibler Divergence (KL).33footnotemark: 3 reflects the acoustic similarity between the generated and reference samples to a certain extent. It is computed over PANNs’ multi-label class predictions.
Density and Coverage [^2] https://github.com/clovaai/generative-evaluation-prdc [60] measures the fidelity and diversity aspects of the generated samples. Fidelity measures how closely the generated samples match the real ones, while diversity assesses whether the generated samples capture the full range of variation found in real samples. We use CLAP[83] embeddings for manifold estimation.
Imagebind Score [^3] https://github.com/facebookresearch/ImageBind [25] assesses to what extent the generated music aligns with the videos. Despite the fact that Imagebind extends the CLIP to six modalities, we only use the branches of audio and vision. Since we use ImageBind to filter out video-audio pairs with a low matching score during dataset construction, the ImageBind score is naturally used in our evaluation. We acknowledge that imageBind is not specifically trained on music data, but there is no better choice currently.
4 Method
Figure 2: Overview of the VidMuse Framework.
In this section, we elaborate on the details of our proposed VidMuse. Our approach leverages Long-Short-Term modeling to generate music that is aligned with video content. Specifically, the Short-Term Module captures localized visual features to ensure the generated music aligns with local variations in the video, while the Long-Term Module models the entire video, providing global context to guide the Short-Term Module in generating more visually consistent music. The experiments in Sec. 5 demonstrate the effectiveness of our proposed method.
4.1 Overview
We illustrate the VidMuse framework’s pipeline in Fig. 2, which converts visual content to music using a structured four-module architecture. The components include (1) a Visual Encoder, (2) a Long-Short-Term Visual Module, (3) a Music Token Decoder, and (4) an Audio Codec Decoder.
Given an input video X, the Visual Encoder extracts feature representations $\bar{\textbf{X}}$. These features are then fed into the Long-Short-Term Visual Module. Specifically, the Short-Term Module captures localized details, while the Long-Term Module models the entire video, providing global context, detailed in Sec. 4.3. This results in the feature representation Z through multi-head cross-attention. The feature Z is then processed by the Music Token Decoder to generate music tokens $\bar{\textbf{Y}}$. Finally, these music tokens are decoded by the Audio Codec to produce the final music $\bar{\textbf{A}}$.
4.2 Visual Encoder
The Visual Encoder processes input video frames, denoted by $\textbf{X}\in\mathbb{R}^{N\times C\times H\times W}$, where $N$ represents the frame count and each frame has dimensions $C\times H\times W$. This component extracts feature representations $\bar{\textbf{X}}\in\mathbb{R}^{N\times P\times D}$ from the video frames. Here, $P$ and $D$ denote the sequence length, including the class token, and the dimensionality of feature vectors, respectively [19, 61, 73, 6].
4.3 Long-Short-Term Visual Module
To capture both local and global visual clues, we model the visual features through the Long-Short-Term Visual Module. Specifically, we evenly samples $N_{l}$ frames from $\bar{\textbf{X}}$ to obtain global feature representations $\bar{\textbf{X}}_{l}\in\mathbb{R}^{N_{l}\times P\times D}$, and extract a continuous segment of $N_{s}$ frames from $\bar{\textbf{X}}$ to form local feature representations $\bar{\textbf{X}}_{s}\in\mathbb{R}^{N_{s}\times P\times D}$. The representations are then refined through Long-term Modeling and Short-term Modeling, respectively, resulting in long-term features $\bar{\textbf{F}}_{l}\in\mathbb{R}^{N_{l}\cdot\text{$N_{H}$}\times D}$ and short-term features $\bar{\textbf{F}}_{s}\in\mathbb{R}^{N_{s}\cdot\text{$N_{H}$}\times D}$, where $N_{H}$ represents the number of attention heads.
LST-Fusion integrates long-term and short-term features by utilizing the cross-attention mechanism in the form of Query ($\mathbf{q}$), Key ($\mathbf{k}$), and Value ($\mathbf{v}$), which is formulated as:
$\mathbf{q}={\bar{\textbf{F}}_{s}}\mathbf{W}_{q},\ \ \ \mathbf{k}=\bar{\textbf{F}}_{l}\mathbf{W}_{k},\ \ \ \mathbf{v}=\bar{\textbf{F}}_{l}\mathbf{W}_{v},$ (1)
$\textbf{Z}^{\prime}=\text{softmax}(\mathbf{q}\mathbf{k}^{T}/\sqrt{C/h})\mathbf{v}+\bar{\textbf{F}}_{s},$ (2)
where $\mathbf{W}_{q}$, $\mathbf{W}_{k}$, $\mathbf{W}_{v}$ $\in\mathbb{R}^{C\times(C/h)}$ denote trainable parameters, with $C$ representing the embedding dimension and $h$ the number of heads. This mechanism allows the model to query global information rather than generating music based solely on local visual features. It guarantees that the generated music is more consistent with the overall video content.
After the cross-attention, a linear projection layer reconfigures $\textbf{Z}^{\prime}$ to align with the Music Token Decoder’s requirements:
$\textbf{Z}=Linear(\textbf{Z}^{\prime}).$ (3)
The $Linear$ that is a fully connected layer projects $\textbf{Z}^{\prime}$ into a space characterized by $\textbf{Z}\in\mathbb{R}^{N_{s}\cdot N_{H}\times M}$, where $M$ corresponds to the intermediate layer dimension within the Music Token Decoder’s transformer architecture.
The Long-Short-Term Visual Module integrates short-term and long-term features of the video and yields a representation Z, ready for the generation of music tokens. The experiment results in Sec. 5 show the efficacy of this module.
4.4 Music Token Decoder
Music and language are both sequential and structured. We thus can try to adapt LLMs to music [76, 34, 15, 4, 11]. To apply LLM techniques to the audio field, several studies have utilized different audio codecs to represent audio information as discrete tokens [90, 40]. These codecs allow audio segments to be encoded into a matrix with dimensions $[K,T_{s}]$, where $K$ is the number of codebooks and $T_{s}$ is the number of sequence timesteps. Some researchers have also investigated various ways to arrange these codebooks [90, 77, 11, 27, 85, 41].
From the previously described steps, we obtain the feature set $\textbf{Z}\in\mathbb{R}^{N_{s}\cdot N_{H}\times M}$, which is then input into the Music Token Decoder’s transformer $\mathcal{G}$ with a linear classification head. This transformation stage produces an output tensor $\bar{\textbf{Y}}\in\mathbb{R}^{K\times S}$ as music tokens, where $K$ is the number of codebooks and $S$ denotes the sequence steps in the musical piece:
$\bar{\textbf{Y}}=Linear(\mathcal{G}(\textbf{Z})),$ (4)
where $\mathcal{G}$ aims to convert video features into a sequence of music features.
4.5 Optimization
Given the codec $\mathcal{C}$, we denote the Audio Encoder as $\mathcal{C}_{\text{encode}}(\cdot)$ and Audio Decoder as $\mathcal{C}_{\text{decode}}(\cdot)$, respectively. The process of encoding the ground truth audio A into its tokenized representation Y can be represented as
$\textbf{Y}=\mathcal{C}_{\text{encode}}(\textbf{A}).$ (5)
The audio decoder $\mathcal{C}_{\text{decode}}(\cdot)$ is utilized solely during the inference stage to decode music tokens into music. In the training stage, an input video X passes through a sequence of operations to generate a predicted representation $\bar{\textbf{Y}}$. With the target representation, Y, is derived by encoding the ground-truth audio A by eq. 5. The objective of our model is to minimize the divergence between $\bar{\textbf{Y}}$ and Y, formalized as the loss function:
$\mathcal{L}=\mathcal{L}(\textbf{Y},\bar{\textbf{Y}}),$ (6)
where $\mathcal{L}$ is cross-entropy loss in practice.
4.6 Inference
When predicting music on videos of arbitrary length, maintaining music consistency and coherence is particularly important. However, it leads to a significant challenge on computational resources due to the quadratic dependency of transformers-based models on sequence length [89, 7]. To address this problem, similar to methods used in previous works [81, 7], we adopt a sliding window approach for inferring the whole video.
During inference, given an input video X with the length of $L$, we set the sliding window length as $L_{s}$ and the length of the window’s overlap as $O$. The value of $t$ is initially set to 00. While $t+L_{s}\leq L$, the inference steps can be summarized as follows: 1) We use the visual encoder to extract feature embeddings $\bar{\textbf{X}}_{l}$ for a given video X. 2) Based on $\bar{\textbf{X}}_{l}$, we collect the feature embeddings within the time window of $[t:t+L_{s}]$ to obtain $\bar{\textbf{X}}_{s}$. 3) Conditioned on $\bar{\textbf{X}}_{l}$ and $\bar{\textbf{X}}_{s}$, the music token decoder predicts the music segment $\bar{\textbf{Y}}$ at the time window of $[t:t+L_{s}-O]$ after the long-short-term visual module. 4) The music tokens $\bar{\textbf{A}}$ are decoded from $\bar{\textbf{Y}}$ by a audio decoder: $\bar{\textbf{A}}=\mathcal{C}_{\text{decode}}(\bar{\textbf{Y}})$. 5) The slide the window moves forward, i.e., $t=t+L_{s}-O$, and we repeat steps 2 to 4 until the end of the video.
5 Experiments
In this section, we first elaborate on the implementation details of our experiment and then present the main results for both our method and the baseline methods using objective metrics and a subjective user study. Then, we demonstrate that our VidMuse exhibits impressive performance on all metrics. Furthermore, we conduct a comprehensive ablation study to explain our design choices and their impact on the performance.
5.1 Implementation details
We choose CLIP [61] as the visual encoder (refer to experiments in Appendix A). Since this work does not focus on audio encoding and decoding, we use Encodec [13] for 32 kHz monophonic audio as our default compression model and use the pretrained transformer model proposed in MusicGen [11]. The training stage utilizes the AdamW optimizer [51] for 22,000 steps with a batch size of 240 samples. The hyperparameters are set to $\beta_{1}=0.9$, $\beta_{2}=0.95$, with a weight decay of 0.1 and gradient clipping at 1.0. A cosine learning rate schedule is employed, incorporating a warm-up phase of 4,000 steps, and an exponential moving average decay of 0.99. The training is conducted on 48 H800 GPUs, with float16 mixed precision. This setup leverages $190$K samples for pre-training and $20$K samples for fine-tuning. A top-k strategy is applied for sampling, retaining the top $250$ tokens with a temperature setting of $1.0$. In the inference stage, we set the sliding window size as $30$s, and the length of the window’s overlap as $5$s.
@@ 5.2 Main Results
Video2Music [37] $1.782$ $144.881$ $18.722$ $0.103$ $0.023$ $0.136$
CMT [16] $1.220$ $85.704$ $8.637$ $0.080$ $0.070$ $0.124$
M2UGen [36] $0.997$ $52.246$ $5.104$ $0.608$ $0.433$ $0.181$
VM-NET [33] $0.899$ $67.480$ $6.252$ $0.986$ $0.383$ $0.147$
VidMuse 0.738 36.171 2.369 1.175 0.710 0.207
We benchmark several state-of-the-art methods, serving as baselines to compare with our method: 1) Caption2Music serving as a naive baseline, which employs the method [2] to extract the video captions and then outputs the music by feeding captions into MusicGen [11]. 2) Video2Music [37] and 3) CMT [16] which botch focus on predicting MIDI notes [64] from videos while our method directly generates music signals. 4) M2UGen [36], a strong baseline, which leverages a language model to connect vision and language, then use MusicGen [11] to generate music from language. 5) VM-NET [33], essentially different from the above methods, retrieves the piece of music from the database for a given video, while other methods predict music by learning from video-music pairs.
In Table 1, VidMuse integrating both global and local visual clues exhibits impressive performance on all metrics. Specifically, compared with Video2Music [37] and CMT [16], VidMuse shows the superiority in the diversity of generated music based on the metrics of density or coverage, which justifies the advantage of directly predicting music signals compared with MIDI notes. In addition, our method even outperforms the strong competitors, i.e., M2UGen. It proves that predicting music directly based on video input can also achieve better performance. Furthermore, compared with a retrieval-based method i.e., VM-NET [33], VidMuse achieves a higher Imagebind score, indicating that the music generated by the learning-based method is more consistent with the video semantics.
@@ 5.3 User Study
In our user study, we randomly sample 600 video-music pairs from the benchmark to conduct an A/B test, which is a widely used subjective evaluation method in the music field [17, 88], across five music generation methods: CMT, M2UGen, Caption2Music, Ground Truth (GT), and VidMuse. The test was distributed among 40 participants, ensuring each method was compared against another 60 times. The evaluation criteria are four-fold: 1) Audio quality: Refers to the sound quality of the audio track; 2) Video-music alignment: Assesses how well the music matches the visual content, e.g., a scene showing a woman crying should ideally be paired with music that sounds sad; 3) Musicality: Measures the aesthetic appeal of music, independent of audio quality, e.g., a piece of music has high audio quality but is out of tune has poor musicality; 4) Overall assessment: A comprehensive evaluation that considers all aspects of the music.
(a) Audio quality.
(b) Video-music alignment.
(a) Musicality.
(b) Overall assessment.
Figure 4: Extended A/B test user study results. Please refer to Sec. 5.3 where we introduce four criteria, i.e., audio quality, video-music alignment, musicality and overall assessment.
Participants are asked to choose the better sample for each criterion. The Fig. 4 illustrates the statistical results of the user study, where the value at matrix[$i$][$j$] ranges from 00 to $100$, indicating the % of times listeners preferred the method correspond to the $i$-row compared to the method correspond to the $j$-column. For example, in Fig. 4 (c), the value of matrix[$2$][$4$] represents that VidMuse outperforms CMT in $77\%$ of cases in terms of Musicality. Across all evaluation criteria, our method surpasses other methods in more than half of the comparisons, except when compared to the ground truth. Overall, these results robustly validate VidMuse’s effectiveness from the subjective evaluation.
@@ 5.4 Ablation Studies
In this section, we conduct comprehensive ablation studies to determine the optimal hyperparameters and design choices. First, we examine the impact of different model design choices on the generation quality. Second, we investigate the influence of the different video sampling rates and the duration of video segments in the Short-Term module on the results. After that, we explore the impact of the model size on performance. More ablation results are in the Appendix A.
@@
Justification of Design Choices. To validate the impact of different model design choices on our generation results and verify the effectiveness of our method, we first design two modules: a short-term modeling module (STM) and a long-term modeling module (LTM). VidMuse-STM aims to ablate the contribution of STM by using only STM without LTM while VidMuse-LTM utilizes only LTM. We present the results in Tab. 2. Based on these results, we learn an insight that local information plays a more importance role in the generation. Second, we implement two variants with Cross-Attention with learnable Queries (CAQ) in our framework. Specifically, CAQ_SL first uses a CAQ where $K$ and $V$ are short-term features and then uses a CAQ where $K$ and $V$ are long-term features. CAQ_LS do it in an opposite order. As shown in Tab. 2, VidMuse outperforms two variants. It seems that the learnable query requires a deeper architecture, as demonstrated by DETR [8] and Q-Former [43], but this manner reduces the model’s efficiency. Furthermore, we also try a baseline using Slowfast [21] and find its performance degenerates in Tab. 2. This may be because the fast path uses a higher frame rate in inputs than the slow path, but their temporal receptive fields are identical. Thus, Slowfast lacks guidance from global clues.
Exploration on model inputs. To explore the impact of different video sampling rates and the duration of video segments in the Short-Term module on performance, we conducted ablation studies on input FPS and short-term segment duration, detailed in Table 3. To intuitively assess the effectiveness of different settings, we employ an Average Rank (AR) metric. The AR metric ranks the results for a metric across all methods within the same table. The ranking result is from $1$ to $N$ (equals to the number of methods within the table), where $1$ is the best and $N$ is the worst. We eventually obtain AR results by averaging the ranking results for all metrics. Note that the AR results cannot be compared across different tables, since this metric is designed to clearly showcase the dominance for each method within one table. From Table 3, we observe that increasing both FPS and duration tends to enhance model capabilities, suggesting that denser frame sampling yields a more detailed video representation, thereby improving music generation. Nevertheless, to balance computational costs and performance, we use a 30-second duration at 2 FPS as our optimal setting.
@@
The impact of model size on performance. For the architecture of the music token decoder, we test three variants with different scales: small ($455$M), medium ($1.9$B), and large ($3.3$B). The medium-sized VidMuse-M is recognized as the default setting if not stated, as it achieves the best results across all metrics, outlined in Table 4. We observe the larger VidMuse-L model fails to deliver proportional improvements. This discrepancy can be partly attributed to limited GPU resources. The model’s performance is near saturation, and the larger model does not exhibit significant advantages. As a result, VidMuse-M is recognized as the optimal choice, achieving a trade-off between performance and efficiency.
@@
Table 3: Ablation studies on Video Duration and FPS.

### v2

@@ 5.2 Evaluation Metrics
To quantitatively evaluate the effectiveness of our model, we employ a series of metrics to assess different models in terms of quality, fidelity, and diversity of the generated music. These metrics include the Frechet Audio Distance (FAD), Frechet Distance (FD), Kullback-Leibler Divergence (KL), as well as Density and Coverage (Naeem et al., 2020). Additionally, we utilize the ImageBind Score (Girdhar et al., 2023) to examine the alignment between the video and the generated music. We acknowledge that ImageBind has limitations as it is not specifically trained on music data, but it currently seems to be a possible option for evaluating the semantic alignment between video and generated music. Further details on metrics can be found in Appendix A.3.
@@ 5.3 Main Results
Video2Music (Kang et al., 2023) $1.782$ $144.881$ $18.722$ $0.103$ $0.023$ $0.136$
CMT (Di et al., 2021) $1.220$ $85.704$ $8.637$ $0.080$ $0.070$ $0.124$
M2UGen (Hussain et al., 2023) $0.997$ $52.246$ $5.104$ $0.608$ $0.433$ $0.181$
VM-NET (Hong et al., 2017) $0.899$ $67.480$ $6.252$ $0.986$ $0.383$ $0.147$
VidMuse 0.734 29.946 2.459 1.250 0.730 0.202
We benchmark several state-of-the-art methods, serving as baselines to compare with our method: 1) Caption2Music a naive baseline that employs the SpaceTimeGPT [^2] https://huggingface.co/Neleac/timesformer-gpt2-video-captioning to extract the video captions and then outputs the music by feeding captions into MusicGen (Copet et al., 2024). 2) Video2Music (Kang et al., 2023) and 3) CMT (Di et al., 2021) which botch focus on predicting MIDI notes (Rothstein, 1995) from videos while our method directly generates music signals. 4) M2UGen (Hussain et al., 2023), a strong baseline, which leverages a language model to connect vision and language, then use MusicGen (Copet et al., 2024) to generate music from language. 5) VM-NET (Hong et al., 2017), essentially different from the above methods, retrieves a piece of music from the database for a given video, while other methods predict music by training on video-music pairs.
In Table 2, VidMuse, with both global and local visual modeling, exhibits impressive performance on all metrics. Specifically, compared with Video2Music (Kang et al., 2023) and CMT (Di et al., 2021), VidMuse shows the superiority in the diversity of generated music based on the density or coverage, which justifies the advantage of directly predicting music signals compared with MIDI notes. In addition, our method even outperforms the strong competitors, i.e., M2UGen. It proves that our method of predicting music directly based on video input can also achieve better performance. Furthermore, compared with a retrieval-based method i.e., VM-NET (Hong et al., 2017), VidMuse achieves a higher Imagebind score, indicating that the music generated by the learning-based strategy is more consistent with the video semantics.
@@ 5.4 User Study
In our user study, we randomly sample 600 video-music pairs from the benchmark to conduct an A/B test, which is a widely used subjective evaluation method in the music field (Donahue et al., 2023; Yuan et al., 2024), across five music generation methods: CMT, M2UGen, Caption2Music, Ground Truth (GT), and VidMuse. The test was distributed among 40 participants, ensuring each method was compared against another 60 times. The evaluation criteria are four-fold: 1) Audio quality: Refers to the sound quality of the audio track; 2) Video-music alignment: Assesses how well the music matches the visual content, e.g., a scene showing a woman crying should ideally be paired with music that sounds sad; 3) Musicality: Evaluates the aesthetic quality of the music, distinct from audio quality. For example, a piece of music may have good audio quality, but if it is out of tune, it would be considered to have poor musicality; 4) Overall assessment: Comprehensively evaluates the performance for models. Participants are asked to choose the better sample for each criterion. The user study is shown in Fig. A1, where the value at matrix[$i$][$j$] ranges from 00 to $100$, indicating the % of times listeners preferred the method in $i$-row compared to the method in $j$-column. For example, in Fig. 5 (c), the value of matrix[$2$][$4$] represents that VidMuse outperforms CMT in $77\%$ of cases in terms of Musicality. Across all criteria, our method surpasses others in more than half of the comparisons, except when compared to the ground truth. Overall, these results thoroughly validate VidMuse’s effectiveness through subjective evaluation.
@@ 5.5 Ablation Studies
In this section, we conduct comprehensive ablation studies, aiming to find the optimal design choices.
@@
Justification of Design Choices. To validate the impact of different model design choices on our generation results and verify the effectiveness of our method, we first design two modules: a short-term modeling module (STM) and a long-term modeling module (LTM). VidMuse-STM aims to ablate the contribution of STM by removing LTM, while VidMuse-LTM utilizes only LTM. Based on the results in Table 3, we gain the insight that local information plays a more important role in the generation. By integrating global guidance with local information, we improve the alignment of the generated music with the overall video content. Second, we implement two variants with Cross-Attention with learnable Queries (CAQ) in our framework. Specifically, CAQ_SL first uses a CAQ where $K$ and $V$ are short-term features and then uses a CAQ where $K$ and $V$ are long-term features. CAQ_LS do it in an opposite order. As shown in Table 3, VidMuse outperforms two variants. It seems that the learnable query requires a deeper architecture, as demonstrated by DETR (Carion et al., 2020) and Q-Former (Li et al., 2023), but this manner reduces the model’s efficiency. Furthermore, we test a baseline using Slowfast (Feichtenhofer et al., 2019) and find its performance degenerates in Table 3. This may be because the fast path uses a higher frame rate in inputs than the slow path, but their temporal receptive fields are identical. Thus, Slowfast still lacks global guidance.
Figure 5: A/B test user study results. We introduce four criteria, i.e., audio quality, video-music alignment, musicality, and overall assessment in Sec. 5.4.
@@
The impact of model size on performance. For the architecture of the music token decoder, we ablate three variants with different scales: small ($455$M), medium ($1.9$B), and large ($3.3$B). The medium-sized VidMuse-M is the default setting unless stated otherwise, as it achieves the best results across all metrics in Table 5. We observe the larger VidMuse-L model fails to deliver proportional improvements. This discrepancy can be partly attributed to limited GPU resources. The model’s performance is near saturation, and the larger model does not exhibit significant advantages. As a result, VidMuse-M is the optimal choice, achieving a trade-off between performance and efficiency.
Visual Encoder. We conduct ablation experiments to study the impact of various visual encoders. As shown in Table 5, we experiment with different visual encoders, including ViT (Dosovitskiy et al., 2020), CLIP (Radford et al., 2021), VideoMAE (Tong et al., 2022), and ViViT (Arnab et al., 2021). Our results show that VidMuse remains robust in processing visual information for music generation across all encoder choices. To balance computational efficiency and generation quality, we select CLIP (Radford et al., 2021) as the default visual encoder if not stated.
More ablation results are provided in Appendix A.1, where we explore the effects of pretraining data scales, different visual encoders and input settings, etc.
@@
Table 4: Ablation studies on model size.
 Models Params Metrics
[…]
  VidMuse-S $455$M $0.854$ $4.695$ $1.349$
VidMuse-M $1.9$B 0.843 2.413 1.487
VidMuse-L $3.3$B $0.873$ $4.041$ $1.330$
Table 5: Ablation studies on video encoders.
     Encoders           Metrics
[…]
  ViViT $0.822$ 37.167 1.433 $451.83$
VideoMAE $0.778$ $37.900$ $1.074$ $360.99$
CLIP 0.753 $38.261$ $1.122$ 141.24
ViT $0.876$ $50.427$ $1.081$ $562.64$
5.6 Qualitative Analysis
In Fig. A2, our qualitative analysis underscores specific limitations for CMT and M2UGen. Specifically, CMT extracts low-level visual cues to generate symbolic music, i.e., MIDI notes. However, its training strategy for symbolic music generation results in discontinuities, particularly for slowly changing or static frames. In these situations, the model fails to predict symbolic music notes, leading to silence. M2UGen utilizes LLMs to fuse multimodal representation and then project LLMs’ embeddings into music via a text-to-music generation model. However, this approach relies on text embeddings as intermediaries, which causes the loss of visual information and restricts the model’s ability to detect nuanced visual variations. As a result, the music generated by this method usually showcases repetitive musical themes and suffers from a lack of diversity. The last row of Fig. A2 demonstrates that our Long-Short-Term (LST) approach is capable of generating music that is rich in diversity and semantically consistent with the video.
6 Conclusion
Limitations. Our work achieves a significant advancement in video-to-music generation, but it still has some limitations. First, the current implementation relies on the EnCodec model (Défossez et al., 2022) trained on a dataset that restricts the system’s capacity to a 32kHz audio sampling rate. This codec exhibits a noticeable reconstruction loss in the audio signal, potentially lowering the quality of the generated music. Second, training large models requires substantial computational resources. Our future work aims to overcome these limitations by integrating advanced codec technologies to enhance audio reconstruction fidelity, and exploring a more efficient method design.
Appendix A Appendix
A.1 Additional Experiments
Table A1: Ablation studies on the ratio of finetuning data.
     Finetuning Data             Metrics
[…]
  0 0.712 $38.184$ $3.956$ $1.125$ $0.583$ $0.181$
10k $0.717$ $34.667$ $2.961$ $0.856$ $0.673$ $0.196$
20k $0.734$ 29.946 2.459 1.250 0.730 0.202
40k $0.776$ $41.075$ $3.557$ $1.094$ $0.726$ $0.195$
60k $0.828$ $40.160$ $2.844$ $0.977$ $0.660$ $0.192$
Exploration on model inputs. To explore the impact of different video sampling rates and the duration of video segments in the Short-Term module on performance, we conducted ablation studies on input FPS and short-term segment duration, detailed in Table A2. To intuitively assess the effectiveness of different settings, we employ an Average Rank (AR) metric. The AR metric ranks the results for a metric across all methods within the same table. The ranking result is from $1$ to $N$ (equals to the number of methods within the table), where $1$ is the best and $N$ is the worst. We eventually obtain AR results by averaging the ranking results for all metrics. Note that the AR results cannot be compared across different tables since this metric is designed to showcase the dominance for each method within one table clearly. From Table A2, we observe that increasing both FPS and duration tends to enhance model capabilities, suggesting that denser frame sampling yields a more detailed video representation, thereby improving music generation. Nevertheless, to balance computational costs and performance, we use a 30-second duration at 2 FPS as our optimal setting.
Codebook Pattern. The exploration of codebook interleaving patterns has attracted attention from researchers across several domains (Zeghidour et al., 2021; Wang et al., 2023; Copet et al., 2024; Yang et al., 2023a; Lan et al., 2023). In our ablation study focusing on the patterns, we find that while the Parallel and Vall-E (Wang et al., 2023) patterns align with the findings for text-to-music generation in MusicGen (Copet et al., 2024), the flattened codebook pattern does not consistently exceed the performance of the delay pattern in tasks of generating music from video. The delay pattern, notable for its relatively low computational cost, is therefore selected for our implementation. The results of this study are presented in table A3.
Table A2: Ablation studies on video duration and FPS.